OpenStudy (anonymous):

The volume of a rectangular box is equal to the length times the width times the height. Write an expression for the volume of a box with length 15 cm, width w cm, and height 10 cm. Be sure to simplify the expression.

OpenStudy (anonymous):

You can begin with the first sentence: The volume of a rectangular box is equal to the length times the width times the height. From that sentence, you can write an expression for the volume of any rectangular box... Volume = length times width times height or using variables (letters) for the quantities, it looks like: V = L * W * H And they give you both length as L = 15 and height as H = 10. They also tell you that you can use "w" for the width. So you just need to re-write that volume expression using 15, w, and 10 in place of L, W, and H.

OpenStudy (anonymous):

okayy whatt???? sorry im not following

OpenStudy (anonymous):

You are supposed to find an expression for the volume. It's like if they said "find the volume", you would just multiply length times width times height, right? It's the same thing here, except you don't know a number for width... you just know it is "w" centimeters.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

So, for example, if they gave you: Length = 15 Width = 2 Height = 10 Then you would find volume to be L * W * H = 15 * 2 * 10 = 300

OpenStudy (anonymous):

But in this case, they don't give you a number for W. It's still just multiplying though: Length = 15 Width = w Height = 10 Volume = 15 * w * 10

OpenStudy (anonymous):

ooooo okay thanks so much you just saved my life!!!!!

OpenStudy (anonymous):

but that's not quite the end... you need to simplify... multiply the 10 and the 15 together...

OpenStudy (anonymous):

150w

OpenStudy (anonymous):

Volume = 15 * 10 * w = 150w Yep, you got it :)

OpenStudy (anonymous):

if you have to include proper units, it is cm^3 or cubic centimeters or whatever format they want.

OpenStudy (anonymous):

yea see i didn't know how to work out the problem to get to that point it was just a bunch of words and numbers to me but yea putting it that way from the original problem always confuses me....thanks tho

OpenStudy (anonymous):

You get better with practice... just train your brain to translate the words to math expressions. It doesn't come all at once, but with practice, you can see how to make a word problem into a math problem. Good luck :)

OpenStudy (anonymous):

aha okay thank you so much have a nice day!:}